OpenStudy (anonymous):

Please need help. Find the center (h,k) and radius r of the circle and then use these to graph the circle. then find any intercepts if any. 4x^2+32x+y^2=0

OpenStudy (anonymous):

i think c(-8,0) and r=8

OpenStudy (anonymous):

thats what i thought as well but as i was looking at an example they had numbers that i had no idea where they came from. how would i graph it?

OpenStudy (raden):

are u sure the equation given like above ? 4x^2+32x+y^2=0 it is not an equation of circle

OpenStudy (anonymous):

@ReadEn , it is it's (2x+8)^2 + (y+0)^2= 64 , 64 is r^2

OpenStudy (raden):

do u know the general formula the equation of circle, @kyusakazaki ?

OpenStudy (anonymous):

@RadEn yeah ,(x-a)^2 + (y-b)^2= r^2, c(a.b)

OpenStudy (raden):

obviously, the coefficient of x^2 and y^2 always be 1 we cant make the coeff of x^2 and y^2 be 1, right ? 4x^2+32x+y^2=0

OpenStudy (raden):

except, it is 4x^2+32x+4y^2=0 :) @channie2003 , is there typo from u ?

OpenStudy (anonymous):

@RadEn so ?

OpenStudy (raden):

still wait confirm from asker, @channie2003 :)

OpenStudy (anonymous):

sorry let me double check im sure it is the right equation

OpenStudy (anonymous):

yeah the equation is exact

OpenStudy (raden):

therefore, the data is wrong :) this site might useful http://www.mathwarehouse.com/geometry/circle/equation-of-a-circle.php

OpenStudy (anonymous):

Thank you guys much :)

OpenStudy (raden):

you're welcome :)

OpenStudy (anonymous):

hey haw are you girl?

OpenStudy (anonymous):

I'm sorry mistake,sorry